OpenStudy (cwrw238):

how about this one 101, 12, 11, 10, ....

Parth (parthkohli):

If it was not for the 101, it'd definitely have been 9.

OpenStudy (cwrw238):

yup

OpenStudy (cwrw238):

hint: its not a 'conventional' series with a formula for nth term

OpenStudy (asnaseer):

I think I know this...

mathslover (mathslover):

is it 5 ? my quess

mathslover (mathslover):

guess *

OpenStudy (asnaseer):

in different bases :)

OpenStudy (asnaseer):

so it should be 1 I believe

OpenStudy (cwrw238):

yes and the next one is 5 also then 5,5,5.......

mathslover (mathslover):

yey .... but I only guessed I will like to hear the right soln

OpenStudy (asnaseer):

101 = 5 in base 2 12 = 5 in base 3 11 = 5 in base 4 10 = 5 in base 5 5 = 5 in base 6

OpenStudy (cwrw238):

i think asnaseer has it

OpenStudy (asnaseer):

typo from me above 5 NOT 1

OpenStudy (cwrw238):

right!

OpenStudy (anonymous):

5,5,5,5,...

hartnn (hartnn):

*nice question, i was thinking of different bases but still could not get this one.

OpenStudy (cwrw238):

jimmyrep gave me this one

OpenStudy (amistre64):

\[\large -\frac{\left(\frac{89(n-1.5)}{|n-1.5|}+\frac{(n-2.5)}{|n-2.5|}+\frac{n-3.5}{|n-3.5|}+5\frac{n-4.5}{|n-4.5|}+106\frac{n-5.5}{|n-5.5|}\right)}{2}\] i can get; ...,5,-101,-101,-101 :)